# A flexible quantum data bus

### v1

@@ Abstract
We have laid the theoretical foundation for routing Bell-states on a 2D cluster state. In the following we discuss three potential use-cases of the parallel zipper-scheme, highlighting its applicability and importance.
@@
The first application corresponds to a long-distance quantum network in which the nodes connect via a 2D cluster state. This setting was e.g. studied in [47, 48], in which the authors presented ways how to create a long-distance cluster state by purifying and merging small buildings blocks. Once the cluster state is established, the goal of the network is to generate Bell-states between any of the nodes in the cluster state on demand, maybe even in parallel. Using the results of this work we find that it is possible to connect any, sufficiently far, sets of communication partners via Bell-states by using the parallel zipper-scheme.
@@
But also centralized communication scenarios with a powerful central unit as studied in [38, 39] benefit from our results. For that purpose, suppose that the nodes of a small scale network, similar to a LAN, connect to a central unit via (potentially multiple) Bell-states. In case the central unit uses internally a local cluster state, and the parties connect via a Bell-state to the border of the cluster state, the central unit is able to establish arbitrary, even parallel, connections amongst the nodes by using our quantum data bus proposal. In other words, the central unit uses the internal cluster to permute connections amongst the (outside) nodes. This allows for a flexible way to operate the central unit, as this unit is capable of connecting any of the connected nodes as necessary and on demand.
@@
The last application we highlight corresponds to integrated quantum devices. In this vision, small quantum devices like sensors connect within the integrated quantum device via a Bell-state to the border of a 2D cluster state, similar in spirit to the previous application. This scenario is inspired by classical computing architectures in which microcontrollers connect through busses in a fixed manner. Our results imply that for quantum computing architectures such a fixed wiring is not necessary, but can be established on-demand in terms of quantum data busses by consuming a cluster state. This introduces flexibility, but also extend-ability to integrated quantum device design.
Conclusion.—
@@
In this letter we have considered pre-established 2D cluster states as resource for quantum networks. We identified the main ingredient to route Bell-states in a cluster state, namely the zipper-scheme. The scheme preserves the cluster state entanglement structure, vital for establishing crossings of Bell-states on a cluster. It further turns out that the zipper-scheme lies at the heart of many building blocks to route Bell-states in a cluster state, like for example ’V’-turns or ’L’-turns. We also demonstrated how to further optimize these building blocks by showing how to utilize them for parallel transport of Bell-states. These building blocks play a key role for routing multiple Bell-states in a long-distance network, but also for small scale networks and the design of integrated quantum processing devices. We also discussed an extension of the diagonal routing to multiparty states like for example GHZ states, and we leave the extension of the other building blocks for future work. As an outlook, it remains an open question how noise in the resource state and the apparatus influences the fidelity of the routed Bell-states. First results for a single target Bell pair [32] indicate robustness against imperfections.
Acknowledgments
@@ Appendix A Graph states
Graph states [45, 46, 27] are a specific class of pure, multipartite entangled quantum states whose correlations correspond to a classical graph $G=(V,E)$. The vertex set $V$ and the edge set $E$ correspond to qubits and binary Ising-type interaction between qubits, respectively. Fig. 3 depicts an example graph state in the center. The figure demonstrates how the graph state transforms under certain local operations, like for example local complementations or Pauli basis measurements. Formally, a graph state |$\ket{G}$ with respect to the graph $G$ is constructed from the set of qubits in the $\ket{+}$ state through the interaction
|$\ket{G}=\prod_{\{a,b\}\in E}C\mathrm{Z}_{ab}\ket{+}^{\otimes|V|},$ (2)
where a controlled-Z gate defined as $diag(1,1,1,-1)$ is applied, if the edge between the qubits $a$ and $b$ exits in the graph $G$. For example the central graph in Fig. 3 comprises four qubits, labeled from one to four. One generates the corresponding graph state by applying controlled-Z gates between the qubits one and two, one and four, one and three, two and three, and two and four, respectively.
Figure 3: In the center is a graph state depicted, whereas vertices correspond to qubits, which are initialized in the $\ket{+}$ state and edges represent maximally entanglement between two qubits. In the corners the resulting graph states are shown, which are caused by local single-qubit Pauli measurements or local complementation as indicated by the labeled arrows.
Alternatively, a graph state |$\ket{G}$ is defined as the common, unique eigenstate with eigenvalue $+1$ of the set of stabilizers
@@ A.1 Local Clifford unitaries and Pauli measurements
Local transformations of graph states play a crucial role in distributed settings. Many of them correspond to graphical manipulations rules on the classical graph $G$, including local, single-qubit Clifford operations and Pauli measurements. For example, local complementation corresponds to applying the unitary operator $\sqrt{K_{a}}$ to the graph state |$\ket{G}$. Graphically, a local complementation on vertex $a$ inverts the sub-graph on the neighbors of $a$. The lower left corner of Fig. 3 demonstrates a local complementation at qubit one, inverting the sub-graph consisting of the qubits two, three and four. A $\sigma_{z}$ measurement on qubit $a$ is equivalent to deleting all edges incident to qubit $a$, and a $\sigma_{y}$ measurement on qubit $a$ corresponds to a local complementation followed by a $\sigma_{z}$ measurement on qubit $a$. Finally, a $\sigma_{x}$ measurement on qubit $a$ is implemented by performing a local complementation on a neighbor $b\in N_{a}$, measuring $a$ in $\sigma_{y}$ basis and performing again a local complementation on $b$. In Fig. 3 possible local Pauli basis measurements on a graph state and their results are summarized.
@@ A.2 State Extraction Methods
It was shown in [49] that it is in general NP-complete to decide if a graph state can be extracted from a given graph state by local Clifford operations, local Pauli measurements and classical communication. Several methods have however been discussed to extract specific target states from resource states in the past. We review two approaches to extract single Bell-states from a given resource state, where we refer to the first one as the isolation strategy and the second one as the X-protocol [31].
@@ A.2.1 Isolation strategy
The idea of the isolation strategy boils down to the isolation of the shortest path between the Bell-state constituents by performing Pauli $Z$ basis measurements on the neighbors. To establish the Bell-state between the end vertices one measures all intermediate qubits along the isolated path using Pauli Y or Pauli X measurements. This strategy has been proposed by Raussendorf et al. [34] to isolate (circuit) wires in a measurement-based quantum computer. This approach comes with several disadvantages. First, it necessitates numerous measurements in the Pauli Z basis to isolate the path, introducing noise to the cluster state. Second, it fails to preserve connectivity (in terms of entanglement) within the cluster, as Pauli Z measurements essentially cut holes into the cluster state. The latter might be problematic, especially if the goal of a protocol corresponds to having multiple, parallel Bell-states.
A.2.2 X-Protocol
@@
An alternative protocol to extract Bell-states from a cluster state has been proposed by Hahn et al. [31], referred to as X-protocol. It first measures all qubits along the shortest path in the Pauli $X$ basis and then removes all qubits still adjacent to the Bell-state by Pauli $Z$ measurements. Ref. [31] shows that the X-protocol requires in the worst case the same number of measurements as the isolation strategy. Note that the butterfly network [50] is the smallest example for the X-protocol. Other works [33, 32] have shown that a staircase-formed measurement path should be preferred over a straight path in a 2D cluster state, because it reduces the remaining Pauli $Z$ basis measurements to the initial neighbors attached to the Bell pair. In the following section we report the potential of the X-protocol for establishing multiplexed communication in a two dimensional cluster state.
Appendix B Proof of Zipper-scheme
@@
In this section we show that the zipper-scheme restores the structure of the remaining cluster state while creating a Bell-state between the two end qubits. We use the graphical and set theoretical rules for Pauli measurements and local Clifford operations from the Refs. [46, 45]. We consider a 2D cluster state where we apply the zipper-scheme along a diagonal path of qubits, labeled from $v_{1}$ to $v_{n}$, to generate a Bell-state between the two end qubits $b_{1}$ and $b_{2}$ of the path, see also Fig. 4. The measurements on the diagonal impact the neighborhood along the path. We distinguish two kinds of qubits in the vicinity of the measured qubits, namely exclusive neighbor qubits that are in the neighborhood of only one measured or the Bell-state qubit, labeled with $e_{i}$ variables, and qubits that are neighbors of at least two measured qubits on the diagonal, labeled with $r_{i}$ variables. The left-most sub-figure in Fig. 4 sketches the situation for a diagonal consisting of six qubits. In the following we explain the zipper-scheme step by step by starting with the Pauli X measurement of $v_{1}$.
Pauli X measurement
@@
Figure 4: Core of zipper-scheme: The aim is to establish the Bell-pair $(b_{1},b_{2})$ with the zipper-scheme on the diagonal $v_{1}$ to $v_{6}$, the initial setting is shown in first step. The second step shows the result of a local complementation on the qubit $b1$, which enlarges the neighborhood of $v_{1}$. The next step shows the Pauli Y measurement on $v_{1}$, which establishes the edge $(b_{1},v_{2})$, necessary for the next measurement, as well as the edge $(r_{1},r_{2})$, necessary for restoring the cluster state, and it removes the qubit $r_{1}$ from the measurement collective, which removes the necessity to measure it in Pauli Z basis. The last step shows the final local complementation on $b_{1}$, which restores a similar configuration as in the initial step, merely $r_{1}$ is replaced by $r_{2}$ and the exclusive neighbors switch from $b_{1}$ to $v_{2}$ and from $v_{1}$ to $b_{1}$.
@@
When measuring $v_{1}$ the first step corresponds to a local complementation on the selected, special neighbor, namely on qubit $b_{1}$. This extends the edge set of the sub-graph induced by the neighbors of $b_{1}$ to a fully-connected sub-graph on the vertex set $\{e_{1},e_{2},r_{1},v_{1}\}$, see the second sub-figure in Fig. 4. Note that this also extends the neighborhood of the measurement qubit $v_{1}$, which we virtually measure in the Pauli Y basis next. We note that this virtual measurement corresponds to a local complementation on $v_{1}$ followed by a Pauli Z measurement. The virtual Pauli Y measurement induces many changes to the edge set of the graph, as one can see in the third sub-figure in Fig. 4. For example, the qubit $r_{1}$ which was a neighbor of two qubits along the measurement path gets detached from both qubits $b_{1}$ and $v_{2}$ on the path. This implies that $r_{1}$ is not a direct neighbor neither to the qubits on the diagonal path nor the Bell-state qubits. Furthermore, this circumstance removes the necessity to measure it in the Pauli Z basis. Also, the virtual Pauli Y measurement creates edges between the qubits $r_{1}$ and $r_{2}$ as well as between $b_{1}$ and $v_{2}$, where the latter is the key ingredient to continue with the X-protocol. Finally, note that exclusive neighbors hop from $b_{1}$ to $v_{2}$ and from $v_{1}$ to $b_{1}$. The final step is another local complementation on the special neighbor $b_{1}$, which removes the fully-connected graph of the vertices $\{v_{2},r_{2},e_{3}\}$ from the graph state, see the last sub-figure in Fig. 4. This last step removes the edge between the next qubit to measure $v_{2}$ and $r_{2}$, and thus, it restores a similar setting as for measuring the qubit $v_{1}$. In this setting we find that $r_{2}$ is a neighbor of the Bell-state qubit $b_{1}$ and $v_{3}$ in a similar manner as $r_{1}$ was for $b_{1}$ and $v_{2}$ in the previous step. Specifically, since $r_{2}$ is now in a similar situation as $r_{1}$ was initially the same observations outlined above apply when measuring $v_{2}$. In summary, when we measure the qubits $v_{1}$ to $v_{n}$, $b_{1}$ connects step-by step to the next qubit until it ends up connecting to $b_{2}$. The exclusive neighbors $e_{i}$ toggle between $b_{1}$ and the qubit to be measured $v_{j}$ in each measurement step $j$. Importantly, all qubits $r_{i}$ detach from the measurement path and connect with each other in such a way that they restore the underlying cluster state structure. The exclusive neighbors $e_{i}$ accompany the Bell states and measurement path until the very end which results in the necessity to remove them in the end.
@@ Appendix C Horizontal and vertical transport of multiple data lines in a 2D cluster state
In this section we show how to transport $n$ parallel Bell-states $(A_{1},B_{1}),\dots,(A_{n},B_{n})$ on a straight horizontal or vertical line in a 2D cluster state. To implement such a parallel transport, we first apply the zipper-scheme to diagonally transport the first data line ($A_{1}$) to the last line ($A_{n}$). By using the cluster state restoring property of the zipper-scheme, we transport in a second step the second data line ($A_{2}$) to the second last position ($A_{n-1}$) again via the zipper-scheme. We repeat this until we reversed the order of all lines, namely to $(A_{n},A_{n-1},\dots A_{1})$. Next, we apply the same measurement strategy again. We observe that we recover the original order of the lines through that. However, we note that we must keep one qubit on each permuted line, resulting in $n$ qubits between the two measurement strategies, to have cluster states before starting the second application of the zipper-scheme. The final step is to apply Pauli Y measurements to merge these intermediate qubits on each line in between to the create the final Bell-states. If we have a distance of $N$ qubits between a Bell pair, we can retrieve from a 2D cluster state at most $\mathcal{O}(N/2)$ Bell-states, and note that a fixed number of measurements, scaling with the number of parallel lines, is required to transport parallel lines. In Fig. 5 we show the measurement strategy to transport four Bell-states in parallel step by step. Notice that in contrast to using the isolation strategy, there are no disconnected lines between data lines. This implies (roughly) a factor of two improvement in required resources to transport multiple data lines in parallel.
Pauli X measurement
Pauli Z measurement
repeat measurement pattern on right-hand side
Pauli Y measurement
@@
Appendix D Taking corners in a 2D cluster state with multiple data lines
In this section we demonstrate how to use the zipper-scheme to change the direction of a quantum data bus from vertical to horizontal (and vice versa). For that purpose, we apply the zipper-scheme (in inward direction) to the outermost data line until we diagonally reach the inner most data line, resulting in a perpendicular turn. We illustrate this step also by the leftmost blue line in Fig. 6. The cluster state preservation property of this first step enables us to apply the zipper-scheme to the second most left qubit until we reach again the innermost qubit. We observe that this results in a perpendicular turn of the second data line of our quantum data bus. Repeatedly applying this pattern results in a full turn of all data lines, as illustrated in Fig. 6.
Pauli X measurement
Pauli Z measurement
Figure 6: The figure shows a step-by-step guide for the measurement pattern to perform a perpendicular turn from horizontal to vertical in a 2D cluster state with four parallel data lines. The zipper-scheme is applied from the left- to right-most qubits, where each application of the zipper-scheme enables the next measurement.
@@ Appendix D Taking corners in a 2D cluster state with multiple data lines
Appendix E Straight to diagonal routing of multiple data lines in a 2D cluster state

### v3

@@ VI Applications
We have laid the theoretical foundation for routing Bell states on a 2D cluster state. In the following we discuss three potential use-cases of the parallel zipper scheme, highlighting its applicability and importance.
@@
The first application corresponds to a long-distance quantum network in which the nodes, which hold a single qubit each, connect via a two-dimensional cluster state. Note that we allow these nodes to be simple clients as well as sophisticated quantum local area networks (LAN) nodes with internal structure. This setting was studied, e.g., in [48, 49], in which the authors presented ways how to create a long-distance cluster state by purifying and merging small buildings blocks. Once the cluster state is established, the goal of the network is to generate Bell states between any of the nodes in the cluster state on demand, maybe even in parallel. Using the results of this work we find that it is possible to connect any, sufficiently far, sets of communication partners via Bell states by using the parallel zipper scheme.
@@
But also centralized communication scenarios with a powerful central unit as studied in [40, 41] benefit from our results. For that purpose, suppose that the nodes of a small scale network, similar to a LAN, connect to a central unit via (potentially multiple) Bell states. In case the central unit uses internally a local cluster state, and the parties connect via a Bell state to the border of the cluster state, the central unit is able to establish arbitrary, even parallel, connections among the nodes by using our quantum data bus proposal. In other words, the central unit uses the internal cluster to permute connections among the (outside) nodes. This allows for a flexible way to operate the central unit, as this unit is capable of connecting any of the connected nodes as necessary and on demand. The advantage of having a pre-established cluster state over distributing Bell pairs is that each node only needs a single quantum memory at each node. Moreover, this method required less latency compared to distributing the Bell state on demand.
@@
The last application we highlight corresponds to integrated quantum devices. In this vision, small quantum devices like sensors connect within the integrated quantum device via a Bell state to the border of a 2D cluster state, similar in spirit to the previous application. This scenario is inspired by classical computing architectures in which microcontrollers connect through buses in a fixed manner. Our results imply that for quantum computing architectures such a fixed wiring is not necessary, but can be established on-demand in terms of quantum data buses by consuming a cluster state. This introduces flexibility, but also extendability to integrated quantum device design.
VII Conclusion
@@
In this paper we have considered pre-established 2D cluster states as resource for quantum networks. We identified the main ingredient to route Bell states in a cluster state, namely the zipper scheme. The scheme preserves the cluster state entanglement structure, vital for establishing crossings of Bell states on a cluster. It further turns out that the zipper scheme lies at the heart of many building blocks to route Bell states in a cluster state, such as, for example 'V'-turns or 'L'-turns. We also demonstrated how to further optimize these building blocks by showing how to utilize them for parallel transport of Bell states. These building blocks play a key role for routing multiple Bell states in a long-distance network, but also for small-scale networks and the design of integrated quantum processing devices. We also discussed an extension of the diagonal routing to multiparty states, such as, for example, GHZ states, and we leave the extension of the other building blocks for future work. As an outlook, it remains an open question how noise in the network state and the apparatus influences the fidelity of the routed Bell states. The first results for a single target Bell pair [34] indicate robustness against imperfections.
VIII Acknowledgments
@@ Appendix A Graph states
[…]
$|,\ket{G}=\prod_{\{a,b\}\in E}CZ_{ab}\ket{+}^{\otimes|V|}$, (2)
where a controlled-$Z$ gate defined as diag(1,1,1,1)diag(1,1,1,-1) is applied, if the edge between the qubits $a$ and $b$ exists in the graph $G$. For example the central graph in Fig. 4 comprises four qubits, labeled from one to four. One generates the corresponding graph state by applying controlled-$Z$ gates between the qubits one and two, one and four, one and three, two and three, and two and four, respectively.
[…]
@@ A.1 Local Clifford unitaries and Pauli measurements
[…]
@@ A.2 State extraction methods
It was shown in [50] that it is in general NP-complete to decide if a graph state can be extracted from a given graph state by local Clifford operations, local Pauli measurements and classical communication. Several methods have however been discussed to extract specific target states from resource states in the past. We review two approaches to extract single Bell states from a given resource state, where we refer to the first one as the isolation strategy and the second one as the X protocol [33].
@@ A.2.1 Isolation strategy
The idea of the isolation strategy boils down to the isolation of the shortest path between the Bell state constituents by performing Pauli $Z$ basis measurements on the neighbors. To establish the Bell state between the end vertices one measures all intermediate qubits along the isolated path using Pauli $Y$ or Pauli $X$ measurements. This strategy has been proposed by Raussendorf et al. [36] to isolate (circuit) wires in a measurement-based quantum computer. This approach comes with several disadvantages. First, it necessitates numerous measurements in the Pauli $Z$ basis to isolate the path, introducing noise to the cluster state. Second, it fails to preserve connectivity (in terms of entanglement) within the cluster, as Pauli $Z$ measurements essentially cut holes into the cluster state. The latter might be problematic, especially if the goal of a protocol corresponds to having multiple, parallel Bell states.
A.2.2 X protocol
@@
An alternative protocol to extract Bell states from a cluster state has been proposed by Hahn et al. [33], referred to as X protocol. It first measures all qubits along the shortest path in the Pauli $X$ basis and then removes all qubits still adjacent to the Bell state by Pauli $Z$ measurements. Ref. [33] shows that the X protocol requires in the worst case the same number of measurements as the isolation strategy. Note that the butterfly network [51] is the smallest example for the X protocol. Other works [35, 34] have shown that a staircase-formed measurement path should be preferred over a straight path in a 2D cluster state, because it reduces the remaining Pauli $Z$ basis measurements to the initial neighbors attached to the Bell pair. In the following section we report the potential of the X protocol for establishing multiplexed communication in a two dimensional cluster state.
Appendix B Proof of zipper scheme
@@
In this section we show that the zipper scheme restores the structure of the remaining cluster state while creating a Bell state between the two end qubits. We use the graphical and set theoretical rules for Pauli measurements and local Clifford operations from the Refs. [47, 46]. We consider a 2D cluster state where we apply the zipper scheme along a diagonal path of qubits, labeled from $v_{1}$ to $v_{n}$, to generate a Bell state between the two end qubits $b_{1}$ and $b_{2}$ of the path, see also Fig. 5. The measurements on the diagonal impact the neighborhood along the path. We distinguish two kinds of qubits in the vicinity of the measured qubits, namely exclusive neighbor qubits that are in the neighborhood of only one measured or the Bell state qubit, labeled with $e_{i}$ variables, and qubits that are neighbors of at least two measured qubits on the diagonal, labeled with $r_{i}$ variables. The left-most sub-figure in Fig. 5 sketches the situation for a diagonal consisting of six qubits. In the following we explain the zipper scheme step by step by starting with the Pauli $X$ measurement of $v_{1}$.
@@
Figure 5: Core of zipper scheme: The aim is to establish the Bell-pair $(b_{1},b_{2})$ with the zipper scheme on the diagonal $v_{1}$ to $v_{6}$, the initial setting is shown in first step. The second step shows the result of a local complementation on the qubit $b1$, which enlarges the neighborhood of $v_{1}$. The next step shows the Pauli $Y$ measurement on $v_{1}$, which establishes the edge $(b_{1},v_{2})$, necessary for the next measurement, as well as the edge $(r_{1},r_{2})$, necessary for restoring the cluster state, and it removes the qubit $r_{1}$ from the measurement collective, which removes the necessity to measure it in Pauli $Z$ basis. The last step shows the final local complementation on $b_{1}$, which restores a similar configuration as in the initial step, merely $r_{1}$ is replaced by $r_{2}$ and the exclusive neighbors switch from $b_{1}$ to $v_{2}$ and from $v_{1}$ to $b_{1}$.
@@
When measuring $v_{1}$ the first step corresponds to a local complementation on the selected, special neighbor, namely on qubit $b_{1}$. This extends the edge set of the subgraph induced by the neighbors of $b_{1}$ to a fully-connected subgraph on the vertex set $\{e_{1},e_{2},r_{1},v_{1}\}$, see the second sub-figure in Fig. 5. Note that this also extends the neighborhood of the measurement qubit $v_{1}$, which we virtually measure in the Pauli $Y$ basis next. We note that this virtual measurement corresponds to a local complementation on $v_{1}$ followed by a Pauli $Z$ measurement. The virtual Pauli $Y$ measurement induces many changes to the edge set of the graph, as one can see in the third sub-figure in Fig. 5. For example, the qubit $r_{1}$ which was a neighbor of two qubits along the measurement path gets detached from both qubits $b_{1}$ and $v_{2}$ on the path. This implies that $r_{1}$ is not a direct neighbor neither to the qubits on the diagonal path nor the Bell state qubits. Furthermore, this circumstance removes the necessity to measure it in the Pauli $Z$ basis. Also, the virtual Pauli $Y$ measurement creates edges between the qubits $r_{1}$ and $r_{2}$ as well as between $b_{1}$ and $v_{2}$, where the latter is the key ingredient to continue with the X protocol. Finally, note that exclusive neighbors hop from $b_{1}$ to $v_{2}$ and from $v_{1}$ to $b_{1}$. The final step is another local complementation on the special neighbor $b_{1}$, which removes the fully-connected graph of the vertices $\{v_{2},r_{2},e_{3}\}$ from the graph state, see the last sub-figure in Fig. 5. This last step removes the edge between the next qubit to measure $v_{2}$ and $r_{2}$, and thus, it restores a similar setting as for measuring the qubit $v_{1}$. In this setting we find that $r_{2}$ is a neighbor of the Bell state qubit $b_{1}$ and $v_{3}$ in a similar manner as $r_{1}$ was for $b_{1}$ and $v_{2}$ in the previous step. Specifically, since $r_{2}$ is now in a similar situation as $r_{1}$ was initially the same observations outlined above apply when measuring $v_{2}$. In summary, when we measure the qubits $v_{1}$ to $v_{n}$, $b_{1}$ connects step-by step to the next qubit until it ends up connecting to $b_{2}$. The exclusive neighbors $e_{i}$ toggle between $b_{1}$ and the qubit to be measured $v_{j}$ in each measurement step $j$. Importantly, all qubits $r_{i}$ detach from the measurement path and connect with each other in such a way that they restore the underlying cluster state structure. The exclusive neighbors $e_{i}$ accompany the Bell states and measurement path until the very end which results in the necessity to remove them in the end.
@@ Appendix C Horizontal and vertical transport of multiple data lines in a 2D cluster state
In this section we show how to transport $n$ parallel Bell states $(A_{1},B_{1}),\dots,(A_{n},B_{n})$ on a straight horizontal or vertical line in a 2D cluster state. To implement such a parallel transport, we first apply the zipper scheme to diagonally transport the first data line ($A_{1}$) to the last line ($A_{n}$). By using the cluster state restoring property of the zipper scheme, we transport in a second step the second data line ($A_{2}$) to the second last position ($A_{n-1}$) again via the zipper scheme. We repeat this until we reversed the order of all lines, namely to $(A_{n},A_{n-1},\dots A_{1})$. Next, we apply the same measurement strategy again. We observe that we recover the original order of the lines through that. However, we note that we must keep one qubit on each permuted line, resulting in $n$ qubits between the two measurement strategies, to have cluster states before starting the second application of the zipper scheme. The final step is to apply Pauli $Y$ measurements to merge these intermediate qubits on each line in between to the create the final Bell states. If we have a distance of $N$ qubits between a Bell pair, we can retrieve from a 2D cluster state at most $\mathcal{O}(N/2)$ Bell states, and note that a fixed number of measurements, scaling with the number of parallel lines, is required to transport parallel lines. In Fig. 6 we show the measurement strategy to transport four Bell states in parallel step by step. Notice that in contrast with using the isolation strategy, there are no disconnected lines between data lines. This implies (roughly) a factor of two improvement in required resources to transport multiple data lines in parallel.
@@
Appendix D Straight to diagonal routing of multiple data lines in a 2D cluster state